OpenStudy (anonymous):

Log(4p-2)=log(-5p+5)

OpenStudy (anonymous):

log cancels out 4p-2=-5p+5 4p+5p=5+2 9p=7 p=7/9

OpenStudy (compassionate):

\(\large Since~you~seem~new~here:\\\bf\Huge{~~\color{red}{\boxed{W}}\color{#FF9200}{\boxed{E}}\color{#FFDB00}{\boxed{L}}\color{#B6ff00}{\boxed{C}}\color{#00ff49}{\boxed{O}}\color{#00DBff}{\boxed{M}}\color{#B600ff}{\boxed{E}}\\~~~~~~~~~~~~~~~~~~\boxed{T}\boxed{O}\\~~~~~~~~~~~\color{#0092ff}{\boxed{O}}\color{#0092ff}{\boxed{P}}\color{#0092ff}{\boxed{E}}\color{#0092ff}{\boxed{N}}\color{#7cc517}{\boxed{S}}\color{#7cc517}{\boxed{T}}\color{#7cc517}{\boxed{U}}\color{#7cc517}{\boxed{D}}\color{#7cc517}{\boxed{Y}}\color{#7cc517}{\boxed{!}}} \\ \large\bf You~can~read~the~\underline{\href{ /code-of-conduct }{Code~of~Conduct}}~here. \\ \it \large Now~let's~try~to~find~an~answer~to~your~problem~together.\)\ \[p = \frac{ 7 }{ 9 }\] \(\bf{\color{green}{-~~OpenStudy~Ambassador:~Compassionate }}\)